# Gemstones: A Model Suite for Multi-Faceted Scaling Laws

### v2

@@ Abstract
Scaling laws are typically fit using a family of models with a narrow range of frozen hyper-parameter choices. In this work we study scaling laws using multiple architectural shapes and hyperparameter choices, highlighting their impact on resulting prescriptions. As a primary artifact of our research, we release the Gemstones: an open-source scaling law dataset, consisting of over 4000 checkpoints from transformers with up to 2 billion parameters and diverse architectural shapes; including ablations over learning rate and cooldown. Our checkpoints enable more complex studies of scaling, such as analyzing the relationship between width and depth. By examining our model suite, we find that the prescriptions of scaling laws can be highly sensitive to the experimental design process and the specific model checkpoints used during fitting.
@@ 3 Designing Our Scaling Laws
To reduce the search space of all possible models, we add some constraints, each of which are either based on precedent from a popular model series like Gemma (Team et al., 2024a, b), Llama (Touvron et al., 2023), Pythia (Biderman et al., 2023), or practical considerations such as hardware details (see Appendix˜A).
@@
For the main set of training runs, we train each model for $350B$ tokens of Dolma 1.7 (Soldaini et al., 2024) data. We target a total batch size of 4~4 million tokens following(Touvron et al., 2023; Dubey et al., 2024; Bai et al., 2023), with a context length of $2048$ and a world batch size of $2048$ sequences. Following Hägele et al. (2024) and Hu et al. (2024), we use a linear learning rate warm up over $80$ million tokens, and then train at a constant learning rate, which we adjust for model size as described in Section˜A.1.
@@ 3.1 Fitting Scaling Laws
We fit scaling laws using methods similar to approach $1$ and $3$ from Chinchilla (Hoffmann et al., 2022). We fit all laws using the log perplexity of all trained models on a sample of $100$ million tokens from a fixed, held-out validation set from the training distribution. We also collect log perplexity values for a range of open source models (Team et al., 2024a, b; Touvron et al., 2023; Dubey et al., 2024; Yang et al., 2024a, b) on the same validation data to allow for a comparison between our predictions and a selection of widely used models. We design a specialized FLOP counting function as we find that simple rules of thumb (e.g., FLOPs per token= $6\times parameters$ (Hoffmann et al., 2022)) do not accurately account for differences in FLOPs between extremely wide and narrow architectures. We discuss this further and present our function in Appendix˜L.
@@
Figure 3: Approach 1 prescriptions. Row one: Validation loss over FLOPs (left) and GPU hours (right) for the first $100$ billion tokens of training. We use Approach 1 to find the optimal points on the convex hull in each setting, marked with black crosses. Row two: We fit a line to the tokens per parameter of empirically optimal models and find a slightly higher, but still constant, tokens per parameter prescription than Hoffmann et al. (2022).  Hoffmann et al. (2022)’s Approach 1 creates $250$ logarithmically-spaced FLOPs bins per order of magnitude, and in red we plot the minimizers over these bins, and the scaling law fitted to these minimizers (binning). Clearly, their Approach 1 is not well-suited for our data, and our convex hull approach is better when we select fewer models to fit our law on. Extended plot in Figure˜19.
Our New Method: The Convex Hull. We fit a lower convex hull to our loss curves. This hull is only supported by a sparse set of optimal models. Fitting on only the vertices of this hull naturally excludes sub-optimal models that lie above the convex hull of optimality, and as we show in Section˜4, this makes the resulting scaling law far more robust to model selection choices.
@@
where $p$ is parameter count and $T$ is tokens. We fit our models using L-BGFS (Liu and Nocedal, 1989) with a Huber loss ($\delta=10^{-4}$) between the empirical log loss and the model prediction, and use multiple initializations following Besiroglu et al. (2024). We ablate to check that our fitting procedure is robust to the size of the grid of initializations and the choice of delta in Section˜K.4.
@@ 4.1 Sizing Up Our Scaling Laws Against Prior Laws and Industry Models
Our convex hull fits the data better than prior approaches.  Hoffmann et al. (2022)’s Approach 1 creates $250$ logarithmically-spaced FLOPs bins per order of magnitude and then uses the models that achieve the best loss in each FLOPs bin to fit the scaling law (a line). However, for our data, their approach does not work very well because it includes many points that are strictly suboptimal with respect to the minimal loss envelope. Our convex hull method omits these points, and fits the line with far fewer points.
In Figure˜3 (row two), we present the prescriptions from our scaling laws for tokens per parameter. We see that the tokens per parameter prescription of our Approach 1 fitting is close to constant, as in Hoffmann et al. (2022), but suggests a slightly larger optimal tokens per parameter ratio than their law. We extend this plot showing predicted total parameters, tokens, and over multiple ablations in Appendix˜K and we give a more detailed plot of each model’s individual validation loss in Appendix˜J. In Appendix˜E, we show a leave-one-out analysis over models when fitting both Approach 1 and 3.
@@ 4.3 Modeling Performance on Different Validation Sets and Downstream Benchmarks
Predicting Benchmark Error
We follow Gadre et al. (2024), predicting downstream average top-1 error (Err) across our benchmarks using the recorded validation loss (L), using a function of the form shown in Equation˜3 where $\epsilon,k,\gamma$ are fit. In Equation˜3, we fit a law to benchmark results sampled at every $10$ billion training tokens, we see that our fitted curve fits the data well. We observe greater variation around the fit compared to Gadre et al. (2024), which we attribute to the considerable differences in the width and depth of the Gemstones models.
Predicting Benchmark Accuracy
Following Bhagia et al. (2024), we calculate task loss by taking the loss over the correct answer to each benchmark question and averaging over all questions. We then use this task loss to predict average task accuracy across $4$ downstream benchmarks. We find the accuracy of ARC, HellaSwag and MMLU to be most predictable at smaller compute scales and use this subset of benchmarks when fitting scaling laws to predict accuracy. Concurrently, Magnusson et al. (2025) also observe this pattern across the same set of benchmarks. We predict average task accuracy by fitting a sigmoidal function of the form shown in Equation˜5 where $a,b,k,l_{0}$ are fit. In Equation˜5, we fit a law to benchmark results sampled at every $10$ billion training tokens. We see a noisy fit and again suspect this is due to the variation in the Gemstones’ width and depth. In Appendix˜D, we hold out the $2$ billion parameter models and show extrapolation for both benchmark scaling laws and Approach 3 loss predictions.
@@
In the previous section we show how deep models appear to achieve better final loss and accuracy on benchmarks when measured in terms of FLOPs. However, another crucial axis for practitioners to consider is the amount of wall time it takes to train a model. In Figure˜7, we visualize in more detail the top of Figure˜3, highlighting in color the approximate aspect ratio of the vertices that form our convex hull when fitting. On the left, we see that deep models are able to achieve a lower loss for a given computational budget (FLOPs) and therefore are selected as the vertices of our convex hull when fitting. However, on the right, we see that when the budget is measured in units of computer time (GPU hours), wider models become more pareto optimal. The concept of “overtraining” is an interesting dimension for further analysis, especially while varying width and depth, as one may be able to overtrain a smaller deep model to reach a low loss value quicker than a larger wider model. We analyze overtraining in greater detail in Appendix˜I.
@@
Finally, we plot the average benchmark accuracy (length normalized) of the Gemstones at $200$, $250$, $300$ and $350$ billion tokens. Figure˜8 shows that the 1B scale models ($1280\times 36$, $2560\times 8$, and $1792\times 18$) yield increasing accuracy with depth when constrained to approximately the same FLOP budget (vertically aligned points). We see similar patterns with the $768\times 45$, $1280\times 15$, $1792\times 7$, and $1024\times 28$ models, as well as the larger 2B models. This result is hinted at in Table 9 of the Gemma 2 report (Team et al., 2024b), where the authors note that for two models at the 9B scale the deeper model slightly outperforms the wider model across downstream benchmarks, but details of the exact experiment are sparse. Recent work suggests deeper layers in networks “do less” than shallower ones and can be pruned away (Gromov et al., 2024), but our downstream evaluations suggest that there are also advantages to additional model depth. We see similar patterns in the individual performance on each benchmark, and include those charts in Appendix Figure˜15.
Figure 8: Benchmark Performance. We benchmark all models using the $200$, $250$, $300$ and $350$ billion token checkpoints. Models show increasing accuracy with depth when constrained to approximately the same FLOP budget (vertically aligned points). This relationship between depth and accuracy can also be observed in many individual benchmarks (Figure˜15).
@@ A.1 Optimal Learning Rates for Gemstones
While $lr_{\text{base}}=5$ achieves stable training for most models under the scheme described above, wider architectures (e.g., $256$ width-depth ratio) occasionally exhibit loss spikes nonetheless. Despite these instabilities, via rollbacks and minor modifications to the learning rates for the most extreme models, all models in the suite are trained to 350350B tokens without divergence. We discuss these issues and our solutions further in Section˜J.2.
@@ Appendix B Extended Related Works
Scaling laws are broadly applied to many areas of machine learning, such as machine translation. Ghorbani et al. [2021] split the parameters term into two, one for each of encoder and decoder, and similarly to Gordon et al. [2021] analyze the relationship between BLEU scores and scaling laws. Zhang et al. [2022] and Bansal et al. [2022] study the impact of architecture choice on the scaling law, finding increasing data or parameters can compensate for worse architectural decisions.
There are multiple works analyzing whether scaling laws can be used to predict downstream performance. Ruan et al. [2024] show that scaling laws can be predictive of benchmark performance. Caballero et al. [2023] observe that traditional scaling laws cannot capture complex behaviors like non-monotonic trends nor inflection points. They propose broken scaling laws–a piecewise or smoothly broken power law–that better predicts performance of both upstream and downstream tasks.
Scaling laws have also been applied to sparse architectures. Clark et al. [2022] analyze how the number of experts can be used in the law, studying both linear and quadratic interactions for many types of routing models. Yun et al. [2024] extend this, analyzing the trade offs between optimal training and optimal inference. Krajewski et al. [2024] find that with optimal settings a Mixture of Experts model always outperforms a transformer model at any computational budget. Frantar et al. [2023] focus on weight sparsity within foundation models, adding a multiplicative parameter on the parameters term in the law. Model shape has also been analyzed for sparse mixture of expert models and in the context of finetuning. Krajewski et al. [2024] use the ratio between the standard feed-forward hidden dimension and the hidden dimension of an individual expert to allow their law for mixture of expert models to predict the width of the experts.
@@
The ability to train a series of models with extremely different parameter counts is an implicit requirement of any scaling law analysis. Work on zero-shot hyperparameter transfer across transformer model widths is mature [Yang et al., 2021, Everett et al., 2024, Hayou and Yang, 2023, Dey et al., 2024]. Achieving transfer across diverse model depths is less well studied, especially in transformer language models [Bordelon et al., 2024, Yang and Hu, 2021, Yang et al., 2023]. While Yang et al. [2023] argue that depth transfer requires scaling in $1/\sqrt{L}$ because this is the unique regime with maximum feature diversity, more recently Dey et al. [2025] argue that instead the scaling should be in $1/L$.
@@ Appendix C Data Sampling
Appendix D Extrapolation to Larger Models
@@
Appendix E Leave-One-model-Out Analysis
@@
Appendix F Variability in fitting
@@
Figure 14: We demonstrate the variability in fitting scaling laws by resampling our data many different ways. We label prescriptions found using Approach 1 with “Approach 11” in the legend, otherwise approach 33 is used. All tokens counts available are used to fit the laws unless stated otherwise in the legend, for example $\leq 100B$ means that only token counts less than or equal to $100B$ are used in fitting.
@@
Appendix G Individual Benchmark Results
In Figure˜15, we see zero-shot MMLU scores of our larger models are quite non-trivial at $28\%-33\%$, despite being trained on an open dataset, without a cooldown period, or any sort of post-training.
Figure 15: Individual benchmark performance. We benchmark all models using the 200, 250, 300, and 350B token checkpoints, converting this to FLOPs using the formula shown in Appendix˜L. Models show increasing accuracy with depth when constrained to approximately the same FLOP budget (vertically aligned points) on many benchmarks. We plot the average benchmark accuracy in Figure˜8.
Appendix H The Price of Stepping Off the Scaling Law

### v3

@@ Abstract
Scaling laws are typically fit using a family of models with a narrow range of frozen hyperparameter choices. In this work we study scaling laws using multiple architectural shapes and hyperparameter choices, highlighting their impact on resulting prescriptions. As a primary artifact of our research, we release the Gemstones: an open-source scaling law dataset, consisting of over 4000 checkpoints from transformers with up to 2 billion parameters and diverse architectural shapes; including ablations over learning rate and cooldown. Our checkpoints enable more complex studies of scaling, such as analyzing the relationship between width and depth. By examining our model suite, we find that the prescriptions of scaling laws can be highly sensitive to the experimental design process and the specific model checkpoints used during fitting.
@@ 3 Designing Our Scaling Laws
To reduce the search space of all possible models, we add some constraints, each of which are either based on precedent from a popular model series like Gemma (Team et al., 2024a, b), Llama (Touvron et al., 2023b), Pythia (Biderman et al., 2023), or practical considerations such as hardware details (see Appendix˜A).
@@
For the main set of training runs, we train each model for $350B$ tokens of Dolma 1.7 (Soldaini et al., 2024) data. We target a total batch size of 4~4 million tokens following(Touvron et al., 2023b; Dubey et al., 2024; Bai et al., 2023), with a context length of $2048$ and a world batch size of $2048$ sequences. Following Hägele et al. (2024) and Hu et al. (2024), we use a linear learning rate warm up over $80$ million tokens, and then train at a constant learning rate, which we adjust for model size as described in Section˜A.1.
@@ 3.1 Fitting Scaling Laws
We fit scaling laws using methods similar to approach $1$ and $3$ from Chinchilla (Hoffmann et al., 2022). We fit all laws using the log perplexity of all trained models on a sample of $100$ million tokens from a fixed, held-out validation set from the training distribution. We also collect log perplexity values for a range of open source models (Team et al., 2024a, b; Touvron et al., 2023b; Dubey et al., 2024; Yang et al., 2024a, b) on the same validation data to allow for a comparison between our predictions and a selection of widely used models. We design a specialized FLOP counting function as we find that simple rules of thumb (e.g., FLOPs per token= $6\times parameters$ (Hoffmann et al., 2022)) do not accurately account for differences in FLOPs between extremely wide and narrow architectures. We discuss this further and present our function in Appendix˜M.
@@
Figure 3: Approach 1 prescriptions. Row one: Validation loss over FLOPs (left) and GPU hours (right) for the first $100$ billion tokens of training. We use Approach 1 to find the optimal points on the convex hull in each setting, marked with black crosses. Row two: We fit a line to the tokens per parameter of empirically optimal models and find a slightly higher, but still constant, tokens per parameter prescription than Hoffmann et al. (2022).  Hoffmann et al. (2022)’s Approach 1 creates $250$ logarithmically-spaced FLOPs bins per order of magnitude, and in red we plot the minimizers over these bins, and the scaling law fitted to these minimizers (binning). Clearly, their Approach 1 is not well-suited for our data, and our convex hull approach is better when we select fewer models to fit our law on. Extended plot in Figure˜20.
Our New Method: The Convex Hull. We fit a lower convex hull to our loss curves. This hull is only supported by a sparse set of optimal models. Fitting on only the vertices of this hull naturally excludes sub-optimal models that lie above the convex hull of optimality, and as we show in Section˜4, this makes the resulting scaling law far more robust to model selection choices. We provide a mathematical definition of our approach in Appendix˜D.
@@
where $p$ is parameter count and $T$ is tokens. We fit our models using L-BGFS (Liu and Nocedal, 1989) with a Huber loss ($\delta=10^{-4}$) between the empirical log loss and the model prediction, and use multiple initializations following Besiroglu et al. (2024). We ablate to check that our fitting procedure is robust to the size of the grid of initializations and the choice of delta in Section˜L.4.
@@ 4.1 Sizing Up Our Scaling Laws Against Prior Laws and Industry Models
Our convex hull fits the data better than prior approaches.  Hoffmann et al. (2022)’s Approach 1 creates $250$ logarithmically-spaced FLOPs bins per order of magnitude and then uses the models that achieve the best loss in each FLOPs bin to fit the scaling law (a line). However, for our data, their approach does not work very well because it includes many points that are strictly suboptimal with respect to the minimal loss envelope. Our convex hull method omits these points, and fits the line with far fewer points. The asymptotic flatness of power law curves makes trying to fit a scaling law an ill-conditioned optimization problem. Our novel convex hull approach is specifically crafted to reduce this variance and our results suggest that when optimal points are sparse, our approach can be used to obtain a more reliable fit (red vs black crosses in Figure˜3)
In Figure˜3 (row two), we present the prescriptions from our scaling laws for tokens per parameter. We see that the tokens per parameter prescription of our Approach 1 fitting is close to constant, as in Hoffmann et al. (2022), but suggests a slightly larger optimal tokens per parameter ratio than their law. We extend this plot showing predicted total parameters, tokens, and over multiple ablations in Appendix˜L and we give a more detailed plot of each model’s individual validation loss in Appendix˜K. In Appendix˜F, we show a leave-one-out analysis over models when fitting both Approach 1 and 3.
@@ 4.3 Modeling Performance on Different Validation Sets and Downstream Benchmarks
Predicting Benchmark Error. We follow Gadre et al. (2024), predicting downstream average top-1 error (Err) across our benchmarks using the recorded validation loss (L), using a function of the form shown in Equation˜3 where $\epsilon,k,\gamma$ are fit. In Equation˜3, we fit a law to benchmark results sampled at every $10$ billion training tokens, we see that our fitted curve fits the data well. We observe greater variation around the fit compared to Gadre et al. (2024), which we attribute to the considerable differences in the width and depth of the Gemstones models.
Predicting Benchmark Accuracy. Following Bhagia et al. (2024), we calculate task loss by taking the loss over the correct answer to each benchmark question and averaging over all questions. We then use this task loss to predict average task accuracy across $4$ downstream benchmarks. We find the accuracy of ARC, HellaSwag and MMLU to be most predictable at smaller compute scales and use this subset of benchmarks when fitting scaling laws to predict accuracy. Concurrently, Magnusson et al. (2025) also observe this pattern across the same set of benchmarks. We predict average task accuracy by fitting a sigmoidal function of the form shown in Equation˜5 where $a,b,k,l_{0}$ are fit. In Equation˜5, we fit a law to benchmark results sampled at every $10$ billion training tokens. We see a noisy fit and again suspect this is due to the variation in the Gemstones’ width and depth. In Appendix˜E, we hold out the $2$ billion parameter models and show extrapolation for both benchmark scaling laws and Approach 3 loss predictions.
@@
In the previous section we show how deep models appear to achieve better final loss and accuracy on benchmarks when measured in terms of FLOPs. However, another crucial axis for practitioners to consider is the amount of wall time it takes to train a model. In Figure˜7, we visualize in more detail the top of Figure˜3, highlighting in color the approximate aspect ratio of the vertices that form our convex hull when fitting. On the left, we see that deep models are able to achieve a lower loss for a given computational budget (FLOPs) and therefore are selected as the vertices of our convex hull when fitting. However, on the right, we see that when the budget is measured in units of computer time (GPU hours), wider models become more pareto optimal. The concept of “overtraining” is an interesting dimension for further analysis, especially while varying width and depth, as one may be able to overtrain a smaller deep model to reach a low loss value quicker than a larger wider model. We analyze overtraining in greater detail in Appendix˜J.
@@
Finally, we plot the average benchmark accuracy (length normalized) of the Gemstones at $200$, $250$, $300$ and $350$ billion tokens. Figure˜8 shows that the 1B scale models ($1280\times 36$, $2560\times 8$, and $1792\times 18$) yield increasing accuracy with depth when constrained to approximately the same FLOP budget (vertically aligned points). We see similar patterns with the $768\times 45$, $1280\times 15$, $1792\times 7$, and $1024\times 28$ models, as well as the larger 2B models. This result is hinted at in Table 9 of the Gemma 2 report (Team et al., 2024b), where the authors note that for two models at the 9B scale the deeper model slightly outperforms the wider model across downstream benchmarks, but details of the exact experiment are sparse. Recent work suggests deeper layers in networks “do less” than shallower ones and can be pruned away (Gromov et al., 2024), but our downstream evaluations suggest that there are also advantages to additional model depth. We see similar patterns in the individual performance on each benchmark, and include those charts in Appendix Figure˜16.
Figure 8: Benchmark Performance. We benchmark all models using the $200$, $250$, $300$ and $350$ billion token checkpoints. Models show increasing accuracy with depth when constrained to approximately the same FLOP budget (vertically aligned points). This relationship between depth and accuracy can also be observed in many individual benchmarks (Figure˜16).
@@ A.1 Optimal Learning Rates for Gemstones
While $lr_{\text{base}}=5$ achieves stable training for most models under the scheme described above, wider architectures (e.g., $256$ width-depth ratio) occasionally exhibit loss spikes nonetheless. Despite these instabilities, via rollbacks and minor modifications to the learning rates for the most extreme models, all models in the suite are trained to 350350B tokens without divergence. We discuss these issues and our solutions further in Section˜K.2.
@@ Appendix B Extended Related Works
Scaling laws have been constructed in many different areas of machine learning since their original proposal in [Kaplan et al., 2020]. Early work on scaling laws for machine translation by Ghorbani et al. [2021] splits the parameters term into two, one for each of encoder and decoder components, and similarly to Gordon et al. [2021] analyzes the relationship between BLEU scores and scaling laws. Subsequently, Zhang et al. [2022] and Bansal et al. [2022] studied the impact of architecture choice on the scaling law, finding increasing data or parameters can compensate for worse architectural decisions. However, the advent of performant open source pipelines for language model development following the release of the Llama series [Touvron et al., 2023a] spurred a renewed flurry of interest in the topic in academic settings.
Architecture
Allen-Zhu and Li [2024] builds scaling laws to model how specific architectural choices impact measures of knowledge acquisition (bits-per-param) in highly controlled settings; they run parallel experiments across dimensions like architecture, quantization, and sparsity to derive insights about which design choices affect acquisition and storage capacity the most. However the more general study of scaling laws for sparse architectures is quite extensive. Clark et al. [2022] analyze how the number of experts can be used in the law, studying both linear and quadratic interactions for many types of routing models. Frantar et al. [2023] focus on weight sparsity within foundation models, adding a multiplicative parameter on the parameters term in the law. Yun et al. [2024] analyzes the trade offs between optimal training and optimal inference and Krajewski et al. [2024] find that with optimal settings, a Mixture of Experts model always outperforms a transformer model at any computational budget. Model shape has also been analyzed for sparse mixture of expert models and in the context of finetuning. Krajewski et al. [2024] use the ratio between the standard feed-forward hidden dimension and the hidden dimension of an individual expert to allow their law for mixture of expert models to predict the width of the experts.
Downstream Benchmarks
Beyond modeling just training loss—the canonical prediction target for scaling laws—there are multiple works analyzing whether scaling laws can be used to predict performance on downstream tasks. Ruan et al. [2024] show that scaling laws can be predictive of benchmark performance. Caballero et al. [2023] observe that traditional scaling laws cannot capture complex behaviors like non-monotonic trends nor inflection points. They propose broken scaling laws–a piecewise or smoothly broken power law–that better predicts performance of both upstream and downstream tasks.
Finetuning and Data-constrained Regimes
Multi-modality
@@
The ability to train a series of models with extremely different parameter counts is an implicit requirement of any scaling law analysis. Bjorck et al. [2024] find that optimal learning rates change with length. Work on zero-shot hyperparameter transfer across transformer model widths is mature [Yang et al., 2021, Everett et al., 2024, Hayou and Yang, 2023, Dey et al., 2024]. Achieving transfer across diverse model depths is less well studied, especially in transformer language models [Bordelon et al., 2024, Yang and Hu, 2021, Yang et al., 2023]. While Yang et al. [2023] argue that depth transfer requires scaling in $1/\sqrt{L}$ because this is the unique regime with maximum feature diversity, more recently Dey et al. [2025] argue that instead the scaling should be in $1/L$.
@@ Appendix C Data Sampling
Appendix D Mathematical Definition of Our Convex Hull Method
We give a mathematical definition for our convex hull method, loosely based on the wikipedia entry for reconstructing functions from epigraphs:
We can define the set of points we have to fit on as FLOPs/GPU hours $(x)$, loss value $(L)$ pairs.
$\mathcal{D}=\{(x_{i},L_{i})\}_{i=1}^{n}\subset\mathbb{R}^{2}$
Let $\text{conv}(\mathcal{D})$ denote the convex hull, the linear interpolation of any two points in $\mathcal{D}$:
$\text{conv}(\mathcal{D})=\left\{\sum_{i=1}^{n}\lambda_{i}(x_{i},L_{i})\ \middle|\ \lambda_{i}\geq 0,\ \sum_{i=1}^{n}\lambda_{i}=1\right\}$
Define
$\hat{L}(x)=\min(\{y|(x,y)\in\mathcal{D}\})$
The lower convex hull is the graph of this new function:
[…]
We think the easiest visualization of this is in Figure˜7 where the red line is the convex hull, the colored crosses are the vertices and the gray lines are all possible points in the dataset.
Appendix E Extrapolation to Larger Models
@@
Appendix F Leave-One-model-Out Analysis
@@
Appendix G Variability in fitting
@@
Figure 14: We demonstrate the variability in fitting scaling laws by resampling our data many different ways. We label prescriptions found using Approach 1 with “Approach 11” in the legend, otherwise approach $3$ is used. All tokens counts available are used to fit the laws unless stated otherwise in the legend, for example $\leq 100B$ means that only token counts less than or equal to $100B$ are used in fitting.
@@
G.1 Variability from sampling
In this section, we visualize the variability in the fitting process due to the frequency of sampling the data. We do this by sampling fitting data every $2$ billion tokens of training and every $10$ billion tokens of training and comparing the laws found. In Figure˜15, comparing the “Ours” lines, we see that for Approach 3 the difference is minimal where as for Approach 1 there is a change in the law. This can be intuitively explained as the data on the hull is much sparser the points the law is fitted to changed, hence for Approach 1 more fitting data gives a more reliable fit. Hence, in all other plots in this paper for Approach 1 we fit with data recorded every $2$ billion tokens for accuracy and for Approach 3 we fit with data every $10$ billion tokens for speed.
Further, in Figure˜15, we also present laws fitted on the DCLM and FineWeb-Edu data in Figure˜4. For the DCLM and FineWeb-Edu data, we record data every $10$ billion tokens of training. In Figure˜15, we see that difference between the laws found by fitting on different data sets compared to our main analysis on Dolma is minimal.
Figure 15: We fit laws when data is sampled every 22 and $10$ billion tokens of training. We also compare to laws fit on DCLM and FineWeb-Edu data. We see sampling frequency is important for Approach 1 and that the difference laws fitted on FineWeb-Edu, DCLM, and Dolma is small.
Appendix H Individual Benchmark Results
In Figure˜16, we see zero-shot MMLU scores of our larger models are quite non-trivial at $28\%-33\%$, despite being trained on an open dataset, without a cooldown period, or any sort of post-training.
Figure 16: Individual benchmark performance. We benchmark all models using the 200, 250, 300, and 350B token checkpoints, converting this to FLOPs using the formula shown in Appendix˜M. Models show increasing accuracy with depth when constrained to approximately the same FLOP budget (vertically aligned points) on many benchmarks. We plot the average benchmark accuracy in Figure˜8.
Appendix I The Price of Stepping Off the Scaling Law